Can you help me solve the equation step by step for the y and x intercepts for y=-3x+1. I think I found the y intercepts but I get confused with the x
the way I solved the first section is: y=-3+1
y: -3(0)+1
y=0+1
y=1 which equals 0,1 y intercepts (?)

Answer by richwmiller(17219)

y intercept is when x=0
y=-3x+1
y= -3(0)+1
y=0+1
y=1 which equals
y intercept is y=1 or (0,1)
x intercept is when y=0
0=-3x+1
3x=1
x intercept =1/3 or (1/3,0)